OpenStudy (anonymous):

How many cubic meters are in a water tank shaped like a cylinder that has a radius of 4 meters and a height of 15 meters?

OpenStudy (watchmath):

\(V==\pi r^2h\) \(V=\pi(4^2)(15)\) \(V=240\pi\quad m^3\)

OpenStudy (anonymous):

thanks